OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of ?

OpenStudy (anonymous):

y = -5x + 22 y = 5x + 22 y=1/5x+32/5 y=-1/5x+32/5 y=-5x+22

OpenStudy (anonymous):

E. y=-1/5x+22

OpenStudy (anonymous):

@kirbykirby @AnswerMyQuestions

OpenStudy (kirbykirby):

There's no number beside your slope in the question.

OpenStudy (anonymous):

-1/5

OpenStudy (kirbykirby):

Your line has the form y = mx + b and m is the slope Here, m = -1/5 Just knowing this only leaves one answer

OpenStudy (anonymous):

2 answers. D and E

OpenStudy (anonymous):

Is the y intercept 22 or 32/5

OpenStudy (kirbykirby):

You could theoretically stop at D, it's the only equation that has a slope of -1/5 (represented in the equation as -1/5x + ... ) If you want to find the slope, you can plug in the x and y value of the point given into the formula and solve for b, the slope: \[7=-\frac{1}{5}(-3)+b \]

OpenStudy (kirbykirby):

oh you corrected your last one E, sorry I didn't realize that

OpenStudy (anonymous):

Yup

OpenStudy (kirbykirby):

then yes in that case, you ca do the above calculation and solve for b

OpenStudy (anonymous):

Did you get E

OpenStudy (kirbykirby):

no

OpenStudy (anonymous):

d

OpenStudy (kirbykirby):

yes, \[ 7=-\frac{1}{5}(-3)+b\\ 7 = \frac{3}{5}+b\\ b = 7 - \frac{3}{5}\\ b = \frac{35}{5}-\frac{3}{5}=\frac{32}{5}\]

OpenStudy (anonymous):

Thanks

OpenStudy (kirbykirby):

yw